Cash Payback Period, Net Present Value Method, and Analysis

Elite Apparel Inc. is considering two investment projects. The estimated net cash flows from each project are as follows:

Year Plant Expansion Retail Store Expansion
1 $168,000 $141,000
2 138,000 165,000
3 119,000 113,000
4 108,000 79,000
5 33,000 68,000
Total $566,000 $566,000

Each project requires an investment of $306,000. A rate of 20% has been selected for the net present value analysis.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162

Required:

1a. Compute the cash payback period for each project.

Cash Payback Period
Plant Expansion
Retail Store Expansion

1b. Compute the net present value. Use the present value of $1 table above. If required, round to the nearest dollar.

Plant Expansion Retail Store Expansion
Present value of net cash flow total $ $
Less amount to be invested $ $
Net present value $ $

2. Because of the timing of the receipt of the net cash flows, the ----------- offers a higher ----------- .
Answer #1

1a)

Plant expansion Retail Store Expansion
Cash flow cummulative cash flow Cash flow cummulative cash flow
0 -306000 --306000 -306000 -306000
1 168000 -306000+168000= -138000 141000 -306000+141000= -165000
2 138000 -138000+138000=0 165000 -165000+165000=0
3 119000 0+119000=119000 113000 113000
4 108000 119000+108000= 227000 79000 192000
5 33000 227000+33000= 260000 68000 260000

payback period = year up to which cummulative cash flow is negative+ [cummulative cash flow of that year /cash flow of next year]

Plant expansion : 1+ [138000/138000]

        = 1+1 =2 years

Retail store expansion = 1+ [165000/165000]

                       = 1+ 1 = 2 years

1b)

Plant expansion Retail store expansion
Present value of net cash flow total

[PVF 20%,1*CF1]+[PVF20%,2*CF2]+[PVF 20%,3*CF3].+[PVF 20%,4*CF4]+[PVF 20%,5*CF5]

[.833*168000]+[.694*138000]+[.579*119000]+[.482*108000]+[.402*33000]

139944+ 95772+ 68901+ 52056+ 13266

369939

[[.833*141000]+[.694*165000]+[.579*113000]+[.482*79000]+[.402*68000]

117453+ 114510+ 65427+ 38078+ 27336

362804

Initial cost (306000) (306000)
NPV 63939 56804

2)Plant expansion

NPV